Find the unknown in the equation log2 48 = log2(2^x+2 minus 2^x)
Please show me the steps to get to the answer, which is x = 4
Stanbon,, you are the man

Answer by stanbon(57361)

log2 48 = log2(2^(x+2) - 2^x)
log2 48 = log2(2^2*2^x - 2^x)
log2 48 = log2(2^x(4-1))
log2 48 = log2(3*2^x)
log2 48 = log2 3 + log2 2^x
log2 48 = log2 3 + x
x = log2 48 - log2 3
x = log2 (48/3)
x = log2 16
x = 4
===============
Cheers,
Stan H.